# Long-Range Graph Wavelet Networks

### v1

@@ 1 Introduction
Long-range interactions are central to complex systems, from electron correlations in quantum chemistry (Ambrosetti et al., 2014; Knörzer et al., 2022) to allosteric effects in biology (Dokholyan, 2016; Zhu et al., 2022). In graph neural networks (GNNs) (Gori et al., 2005; Scarselli et al., 2009; Gilmer et al., 2017; Bronstein et al., 2017), capturing these interactions requires models that can propagate information beyond local neighborhoods without the computational overhead of dense global interactions (Alon & Yahav, 2021; Dwivedi et al., 2022).
Wavelet-based graph neural networks (WGNNs) (Hammond et al., 2011), inspired by wavelet theory (Mallat, 1989), provide a principled framework for this challenge, defining spectral filters that can in principle capture different scales of information propagation through their filtering characteristics. However, designing exact wavelet filters face a fundamental computational bottleneck, as it would require computing the full eigenvalue decomposition of the graph Laplacian, which is prohibitive for large graphs.
To circumvent this cost, current WGNNs (Hammond et al., 2011; Xu et al., 2019; Liu et al., 2024a) approximate wavelet filters using low-order polynomials. While computationally efficient, this approach suffers from two critical limitations. First, polynomial filters aggregate information only within a finite number of hops (Balcilar et al., 2020), inherently restricting their spatial reach. More fundamentally, polynomial approximations face an inherent trade-off between computational efficiency and functional expressiveness: while high-degree polynomials can theoretically approximate any function defined on the real interval (Weierstrass theorem), practically feasible orders cannot accurately represent the discontinuous or steep characteristics needed for selective frequency filtering, such as wavelets (Geisler et al., 2024).
@@
We propose Long-Range Graph Wavelet Networks (LR-GWN), which overcome the limitations of polynomial-based WGNNs through a hybrid filter design. Our method decomposes each wavelet filter into two components: a low-order polynomial for efficient local aggregation and a spectral filter operating on the low-frequency eigenspace to enable long-range propagation (Fig.˜1d). This approach requires only a partial eigendecomposition at preprocessing time, making it computationally practical while achieving the global reach that purely polynomial methods cannot provide. LR-GWN can operate under strict wavelet theory or with relaxed constraints for enhanced performance, providing a principled yet flexible framework for wavelet-based long-range graph learning.
@@ 2 Background
Graphs. We consider undirected graphs $\smash{{\mathcal{G}}=({\bm{A}},{\bm{X}})}$ with adjacency matrix $\smash{{\bm{A}}\in\{0,1\}^{n\times n}}$ and node features $\smash{{\bm{X}}\in\mathbb{R}^{n\times d}}$. Throughout we use the symmetrically normalized Laplacian $\smash{{\mathcal{L}}={\bm{I}}-{\bm{D}}^{-\frac{1}{2}}{\bm{A}}{\bm{D}}^{-\frac{1}{2}}}$, with $\smash{{\bm{D}}=\mathrm{diag}({\bm{A}}\bm{1})}$, which admits eigendecomposition $\smash{{\mathcal{L}}={\bm{U}}{\bm{\Lambda}}{\bm{U}}^{\top}}$ with orthogonal eigenvectors ${\bm{U}}\in\mathbb{R}^{n\times n}$ and eigenvalues $\smash{{\bm{\Lambda}}=\mathrm{diag}(\lambda_{1},\dots,\lambda_{n})}\in\mathbb{R}^{n\times n}$, $\smash{0=\lambda_{1}\leq\dots\leq\lambda_{n}\leq 2}$ (Chung, 1997). The eigenvectors $\smash{{\mathcal{L}}} fo$rm an orthogonal basis, often called the graph Fourier basis. Any graph signal $\smash{{\bm{x}}\in\mathbb{R}^{n}}$ can be decomposed in this basis via the graph Fourier transform (GFT) as $\smash{\widehat{{\bm{x}}}={\bm{U}}^{\top}{\bm{x}}}$, with inverse GFT $\smash{{\bm{x}}={\bm{U}}\widehat{{\bm{x}}}}$. Spectral filtering applies a kernel $\smash{g:\mathbb{R}\to\mathbb{R}}$ by modulating its Fourier coefficients: $\smash{g\ast_{{\mathcal{G}}}{\bm{x}}={\bm{U}}\,\widehat{g}({\bm{\Lambda}})\,{\bm{U}}^{\top}{\bm{x}}}$, where $\smash{\widehat{g}({\bm{\Lambda}})=\mathrm{diag}(\widehat{g}(\lambda_{1}),\dots,\widehat{g}(\lambda_{n}))}$. Small eigenvalues correspond to smooth/global variations, large ones to oscillatory/local patterns (Kreuzer et al., 2021, see Fig. 3 for an intuitive visualization).
Wavelets. Wavelet transforms extend classic Fourier analysis by providing joint localization in both frequency and space (Mallat, 1989, 2009). A mother wavelet $\psi$ generates dilated and translated functions $\smash{\psi_{s,l}(t)=s^{-1/2}\psi\!\left(\tfrac{t-l}{s}\right)}$, whose coefficients $\smash{W_{f}(s,l)=\int_{\mathbb{R}}f(t)\psi^{*}_{s,l}(t)dt}$ decompose a signal $f$ across scales $s$ and locations $l$. Reconstruction requires the admissibility condition, $\smash{\widehat{\psi}(0)=0}$, ensuring $\smash{\widehat{\psi}}$ acts as a band-pass filter. A complementary scaling function $\phi$ with $\smash{\widehat{\phi}(0)>0}$ provides low-pass coverage, yielding a multiresolution analysis where wavelets capture localized variations and the scaling function captures global trends.
@@ 3 Limitations of Polynomial Graph Filters
Perspective I: Propagation. Balcilar et al. (2020) show that an $l$-layer message-passing GNN corresponds exactly to an $l$-order polynomial filter. In message-passing networks, each layer typically uses a first-order update of the form $\smash{{\bm{H}}^{(l)}=(\omega_{0}{\bm{I}}+\omega_{1}{\mathcal{L}}){\bm{H}}^{(l-1)}}$, where each application of $\mathcal{L}$} aggregates information from one-hop neighbors (Kipf & Welling, 2016). Stacking $l$ such layers yields a polynomial filter of order $l$, enabling information flow across at most $l$ hops. Beyond this range, information is inaccessible, and increasing depth triggers over-squashing (Alon & Yahav, 2021), where distant signals collapse into compressed representations. Thus polynomial filters cannot achieve effective long-range propagation without prohibitive computational costs or architectural limitations.
Perspective II: Approximation. Polynomials of low degree are also weak function approximators. By Markov’s inequality (Markov, 1889), the slope of a degree-$\rho$ polynomial on [1,1][-1,1] is bounded as $\max_{x\in[-1,1]}\big{|}P^{\prime}(x)\big{|}\leq\rho^{2},$ which restricts the ability of the polynomial to approximate functions with sharp and steep transitions. Smooth functions can be represented with modest degree, but scaling functions and low-scale wavelets, which exhibit steep cutoffs, require prohibitively high $\rho$. This confines polynomial filters to overly smooth spectral responses.
Perspective III: Convergence. The most fundamental limitation arises with discontinuous filters such as ideal band-pass or virtual-node mechanisms. Geisler et al. (2024) show that polynomial approximations converge arbitrarily slowly in operator norm, regardless of degree. In other words, there exists no sequence of polynomials that can efficiently approximate such filters across graphs. This establishes a structural gap between polynomial parametrizations and the expressive, frequency-selective operators required for modeling long-range dependencies.
@@ 4 Long-Range Graph Wavelet Networks
We propose the Long-Range Graph Wavelet Network (LR-GWN), a parametrization of graph wavelet filters that enables efficient, principled, and expressive filtering of graph signals, specifically tailored to capture long-range dependencies. Our design addresses three core limitations of prior approaches: it resolves the locality bottleneck of polynomial filters, preserves the interpretability of pure wavelet propagation, and achieves linear complexity with respect to the number of edges. The following subsections introduce its hybrid filter parametrization, wavelet-based propagation mechanism, and efficient implementation.
@@ 4.1 Hybrid Parametrization of Wavelet Filters
Truncated spectrum. In practice, we compute a partial eigendecomposition $({\bm{U}},{\bm{\Lambda}})=\operatorname{EVD}({\mathcal{L}},k)$ of lowest $k$ eigenpairs with $\smash{k\ll n}$, i.e., ${\bm{U}}\in\mathbb{R}^{n\times k},{\bm{\Lambda}}\in\mathbb{R}^{k\times k}$. This focuses the spectral component on informative frequencies (e.g., low $\lambda$ for long-range effects) while maintaining scalability. The decomposition costs ${\mathcal{O}}(km)$, where $m\ll n^{2}$ is the number of edges, is computed once at preprocessing time and reused across layers.
Filter parametrization. We implement the spatial filter $P({\mathcal{L}})$ using a Chebyshev basis for stable, recursive evaluation, and the spectral filter $S({\bm{\Lambda}})$ via Gaussian smearing (Schütt et al., 2017), enabling fine-grained, band-specific control. Implementation details are given in Section˜D.1.
@@ 4.2 Principled Wavelet Propagation
We can ensure the band-pass behavior of the wavelet family, hence its wavelet theoretical validity, by enforcing the admissibility condition $\smash{\widehat{\psi}(0)=0}$, which guarantees invertibility of the wavelet transform. In practice, strict admissibility is not always necessary and relaxing it can yield slight empirical gains. A distinctive feature of LR-GWN is that it supports both regimes: admissibility can be enforced when theoretical guarantees are required, or relaxed when flexibility and performance are prioritized. We show how to enforce the constraint in Section˜D.2.2.
@@ 4.3 Computational Efficiency
Summary.
LR-GWN offers a hybrid parametrization that unifies polynomial and spectral filters, a propagation scheme based solely on wavelet operators, and an implementation that scales linearly with the number of edges, as common GNNs. Together, these components provide an efficient, principled, and expressive framework for long-range graph signal filtering.
@@ 5 Empirical Results
Experimental setup. We conduct between 5 and 10 independent runs per experiment using different random seeds and report the mean performance and standard error of the mean. Our experiments are optimized for resource efficiency, requiring less than 10 GB of GPU memory, making them feasible on widely available hardware such as Nvidia GTX 1080Ti and 2080Ti GPUs. We use the AdamW optimizer (Loshchilov & Hutter, 2018) with cosine annealing scheduler (Loshchilov & Hutter, 2017) and linear warmup for stable training. Early stopping based on validation metrics prevents overfitting. All results use the independent filter parametrization (see Section˜D.1). We follow standard practice by augmenting node features with positional encodings (PEs) (Tönshoff et al., 2023), noting that our partial EVD provides PEs for free. Additional implementation details are provided in Appendix˜D.
Wavelet methods categorization. To fairly assess wavelet-based approaches, we distinguish between methods based on their theoretical adherence. Theory-compliant (TC) methods satisfy admissibility constraints and propagate entirely through wavelet operations. Only SGWT (Hammond et al., 2011) and LR-GWN (with admissibility) qualify as theory-compliant. Theory-relaxed (TR) methods either incorporate auxiliary non-wavelet components (WaveGC (Liu et al., 2024a)) or relax theoretical constraints (LR-GWN (without admissibility), GWNN (Xu et al., 2019), DEFT (Bastos et al., 2023), ASWT-SGNN (Liu et al., 2024b)).
@@ 5.1 Long-Range Tasks
Table 1: Long-range benchmark results by theoretical adherence. Theory-compliant methods (TC) satisfy wavelet admissibility and use pure wavelet propagation. Theory-relaxed methods (TR) either use additional non-wavelet modules, or relax theory constraints. Bold/underline represent first and second best, respectively.
@@
SGWT (Hammond et al., 2011) 60.23 $\pm$ 0.27 - 25.39 $\pm$ 0.21 -
GWNN (Xu et al., 2019) - 65.47 $\pm$ 0.48 - 27.34 $\pm$ 0.04
DEFT (Bastos et al., 2023) - 66.95 $\pm$ 0.63 - 25.06 $\pm$ 0.13
WaveGC (Liu et al., 2024a) - 69.73 $\pm$ 0.43 - 24.95 $\pm$ 0.07
@@
Datasets. We evaluate on Peptides-Func (multi-label classification of 10 functional classes) and Peptides-Struct (node-level regression for 11 3D structural properties) from the Long-Range Graph Benchmark (Dwivedi et al., 2023; Tönshoff et al., 2023). These tasks specifically test long-range propagation in biological structures where distant interactions determine biological functions.
LR-GWN achieves state-of-the-art performance on both datasets, as reported in Table˜1. Most notably, it significantly outperforms all existing theory-compliant methods (70.52 vs 60.23 on Peptides-Func) while maintaining theoretical guarantees. When constraints are relaxed, it surpasses all theory-relaxed baselines. It is worth noting that even the theory-compliant version of LR-GWN, achieves comparable performance to the theory-relaxed baselines, while maintaining pure wavelet interpretability and propagation.
5.2 Short-Range Tasks
Datasets. We evaluate on Photo and Computers from Amazon co-purchase graphs (McAuley et al., 2015; Shchur et al., 2018). These datasets focus on local neighborhood interactions, validating that our long-range optimization does not compromise local modeling capabilities.
Table 2: Short-range benchmark results by theoretical adherence. Theory-compliant methods (TC) satisfy wavelet admissibility and use pure wavelet propagation. Theory-relaxed methods (TR) either use additional non-wavelet modules, or relax theory constraints. Bold/underline represent first and second best, respectively.
@@ 5.2 Short-Range Tasks
SGWT (Hammond et al., 2011) 92.45 $\pm$ 0.15 - 85.19 $\pm$ 0.22 -
GWNN (Xu et al., 2019) - 94.45 $\pm$ 0.18 - 90.75 $\pm$ 0.16
ASWT-SGNN (Liu et al., 2025) - 93.80 $\pm$ 0.12 - 89.40 $\pm$ 0.19
WaveGC (Liu et al., 2024a) - 95.37 $\pm$ 0.44 - 92.26 $\pm$ 0.18
@@
As reported in Table˜2, LR-GWN achieves state-of-the-art performance on Photo and competitive results on Computers, demonstrating our proposed hybrid parametrization offers the flexibility to model both global and local propagation. Contrary to long-range tasks, relaxing admissibility constraints does not consistently improve performance on these local interaction datasets.
Training Time. On Photo, training LR-GWN takes approximately 3 seconds per epoch with total training time around 20 minutes for the full run (400 epochs). The partial EVD preprocessing with $k=50$ costs roughly 1.3 seconds using sparse solvers, accounting for approximately 0.1% of total training time.

### v3

# New Perspectives in Advancing Graph Machine Learning

@@ 1 Introduction
Long-range interactions are central to complex systems, from electron correlations in quantum chemistry [2, 25] to allosteric effects in biology [12, 48]. In graph neural networks (GNNs) [21, 40, 19, 8], capturing these interactions requires models that can propagate information beyond local neighborhoods without the computational overhead of dense global interactions [1, 14].
Wavelet-based graph neural networks (WGNNs) [22], inspired by wavelet theory [33], provide a principled framework for this challenge, defining spectral filters that can in principle capture different scales of information propagation through their filtering characteristics. However, designing exact wavelet filters face a fundamental computational bottleneck, as it would require computing the full eigenvalue decomposition of the graph Laplacian, which is prohibitive for large graphs.
To circumvent this cost, current WGNNs [22, 47, 28] approximate wavelet filters using low-order polynomials. While computationally efficient, this approach suffers from two critical limitations. First, polynomial filters aggregate information only within a finite number of hops [4], inherently restricting their spatial reach. More fundamentally, polynomial approximations face an inherent trade-off between computational efficiency and functional expressiveness: while high-degree polynomials can theoretically approximate any function defined on the real interval (Weierstrass theorem), practically feasible orders cannot accurately represent the discontinuous or steep characteristics needed for selective frequency filtering, such as wavelets [18].
@@
We propose Long-Range Graph Wavelet Networks (LR-GWN), which overcome the limitations of polynomial-based WGNNs through a hybrid filter design. Our method decomposes each wavelet filter into two components: a low-order polynomial for efficient local aggregation and a spectral filter operating on the low-frequency eigenspace to enable long-range propagation (Fig.˜1d). This approach requires only a partial eigendecomposition at preprocessing time, making it computationally practical while achieving the global reach that purely polynomial methods cannot provide.
LR-GWN can operate under strict wavelet theory or with relaxed constraints for enhanced performance, providing a principled yet flexible framework for wavelet-based long-range graph learning.
@@ 2 Background
Graphs. We consider undirected graphs $\smash{{\mathcal{G}}=({\bm{A}},{\bm{X}})}$ with adjacency matrix $\smash{{\bm{A}}\in\{0,1\}^{n\times n}}$ and node features $\smash{{\bm{X}}\in\mathbb{R}^{n\times d}}$. Throughout we use the symmetrically normalized Laplacian $\smash{{\mathcal{L}}={\bm{I}}-{\bm{D}}^{-\frac{1}{2}}{\bm{A}}{\bm{D}}^{-\frac{1}{2}}}$, with $\smash{{\bm{D}}=\mathrm{diag}({\bm{A}}\bm{1})}$, which admits eigendecomposition $\smash{{\mathcal{L}}={\bm{U}}{\bm{\Lambda}}{\bm{U}}^{\top}}$ with orthogonal eigenvectors ${\bm{U}}\in\mathbb{R}^{n\times n}$ and eigenvalues $\smash{{\bm{\Lambda}}=\mathrm{diag}(\lambda_{1},\dots,\lambda_{n})}\in\mathbb{R}^{n\times n}$, $\smash{0=\lambda_{1}\leq\dots\leq\lambda_{n}\leq 2}$ [10]. The eigenvectors $\smash{{\mathcal{L}}} fo$rm an orthogonal basis, often called the graph Fourier basis. Any graph signal $\smash{{\bm{x}}\in\mathbb{R}^{n}}$ can be decomposed in this basis via the graph Fourier transform (GFT) as $\smash{\widehat{{\bm{x}}}={\bm{U}}^{\top}{\bm{x}}}$, with inverse GFT $\smash{{\bm{x}}={\bm{U}}\widehat{{\bm{x}}}}$. Spectral filtering applies a kernel $\smash{g:\mathbb{R}\to\mathbb{R}}$ by modulating its Fourier coefficients: $\smash{g\ast_{{\mathcal{G}}}{\bm{x}}={\bm{U}}\,\widehat{g}({\bm{\Lambda}})\,{\bm{U}}^{\top}{\bm{x}}}$, where $\smash{\widehat{g}({\bm{\Lambda}})=\mathrm{diag}(\widehat{g}(\lambda_{1}),\dots,\widehat{g}(\lambda_{n}))}$. Small eigenvalues correspond to smooth/global variations, large ones to oscillatory/local patterns [26, see Fig. 3 for an intuitive visualization].
Wavelets. Wavelet transforms extend classic Fourier analysis by providing joint localization in both frequency and space [33, 34]. A mother wavelet $\psi$ generates dilated and translated functions $\smash{\psi_{s,l}(t)=s^{-1/2}\psi\!\left(\tfrac{t-l}{s}\right)}$, whose coefficients $\smash{W_{f}(s,l)=\int_{\mathbb{R}}f(t)\psi^{*}_{s,l}(t)dt}$ decompose a signal $f$ across scales $s$ and locations $l$. Reconstruction requires the admissibility condition, $\smash{\widehat{\psi}(0)=0}$, ensuring $\smash{\widehat{\psi}}$ acts as a band-pass filter. A complementary scaling function $\phi$ with $\smash{\widehat{\phi}(0)>0}$ provides low-pass coverage, yielding a multiresolution analysis where wavelets capture localized variations and the scaling function captures global trends.
@@ 3 Limitations of Polynomial Graph Filters
Perspective I: Propagation. Balcilar et al. [4] show that an $l$-layer message-passing GNN corresponds exactly to an $l$-order polynomial filter. In message-passing networks, each layer typically uses a first-order update of the form $\smash{{\bm{H}}^{(l)}=(\omega_{0}{\bm{I}}+\omega_{1}{\mathcal{L}}){\bm{H}}^{(l-1)}}$, where each application of $\mathcal{L}$} aggregates information from one-hop neighbors [24]. Stacking $l$ such layers yields a polynomial filter of order $l$, enabling information flow across at most $l$ hops. Beyond this range, information is inaccessible, and increasing depth triggers over-squashing [1], where distant signals collapse into compressed representations. Thus polynomial filters cannot achieve effective long-range propagation without prohibitive computational costs or architectural limitations.
Perspective II: Approximation. Polynomials of low degree are also weak function approximators. By Markov’s inequality [35], the slope of a degree-$\rho$ polynomial on [1,1][-1,1] is bounded as $\max_{x\in[-1,1]}\big|P^{\prime}(x)\big|\leq\rho^{2},$ which restricts the ability of the polynomial to approximate functions with sharp and steep transitions. Smooth functions can be represented with modest degree, but scaling functions and low-scale wavelets, which exhibit steep cutoffs, require prohibitively high $\rho$. This confines polynomial filters to overly smooth spectral responses.
Perspective III: Convergence. The most fundamental limitation arises with discontinuous filters such as ideal band-pass or virtual-node mechanisms. Geisler et al. [18] show that polynomial approximations converge arbitrarily slowly in operator norm, regardless of degree. In other words, there exists no sequence of polynomials that can efficiently approximate such filters across graphs. This establishes a structural gap between polynomial parametrizations and the expressive, frequency-selective operators required for modeling long-range dependencies.
@@ 4 Long-Range Graph Wavelet Networks
We propose the Long-Range Graph Wavelet Network (LR-GWN), a parametrization of graph wavelet filters that enables efficient, principled, and expressive filtering of graph signals, specifically tailored to capture long-range dependencies. Our design addresses three core limitations of prior approaches: it resolves the locality bottleneck of polynomial filters, preserves the interpretability of pure wavelet propagation, and achieves linear complexity with respect to the number of edges.
@@ 4.1 Hybrid Parametrization of Wavelet Filters
Truncated spectrum. In practice, we compute a partial eigendecomposition $({\bm{U}},{\bm{\Lambda}})=\operatorname{EVD}({\mathcal{L}},k)$ of lowest $k$ eigenpairs with $\smash{k\ll n}$, i.e., ${\bm{U}}\in\mathbb{R}^{n\times k},{\bm{\Lambda}}\in\mathbb{R}^{k\times k}$. This focuses the spectral component on informative frequencies (e.g., low $\lambda$ for long-range effects) while maintaining scalability. In practice, truncating to the lowest $k$ eigenpairs preserves most of the signal power, since long-range interactions are dominated by low-frequency modes. The decomposition costs ${\mathcal{O}}(km)$, where $m\ll n^{2}$ is the number of edges, is computed once at preprocessing time and reused across layers.
Filter parametrization. We implement the spatial filter $P({\mathcal{L}})$ using a Chebyshev basis for stable, recursive evaluation, and the spectral filter $S({\bm{\Lambda}})$ via Gaussian smearing [41], enabling fine-grained, band-specific control. Implementation details are given in Section˜D.1.
@@ 4.2 Principled Wavelet Propagation
We can ensure the band-pass behavior of the wavelet family, hence its wavelet theoretical validity, by enforcing the admissibility condition $\smash{\widehat{\psi}(0)=0}$, which guarantees invertibility of the wavelet transform. In practice, strict admissibility is not always necessary and relaxing it can yield slight empirical gains. A distinctive feature of LR-GWN is that it supports both regimes: admissibility can be enforced when theoretical guarantees are required, or relaxed when flexibility and performance are prioritized. In this context, the theoretical guarantees refer to the invertibility of the transform and the preservation of band-pass behavior under the admissibility condition $\smash{\widehat{\psi}(0)=0}$. Enforcing admissibility ensures that LR-GWN defines a valid wavelet transform with full reconstruction capability, while the relaxed variant trades strict invertibility for small empirical gains. We show how to enforce the constraint in Section˜D.2.2.
@@ 5 Empirical Results
Experimental setup. We conduct between 5 and 10 independent runs per experiment using different random seeds and report the mean performance and standard error of the mean. Our experiments are optimized for resource efficiency, requiring less than 10 GB of GPU memory, making them feasible on widely available hardware such as Nvidia GTX 1080Ti and 2080Ti GPUs. We use the AdamW optimizer [32] with cosine annealing scheduler [31] and linear warmup for stable training. Early stopping based on validation metrics prevents overfitting. All results use the independent filter parametrization (see Section˜D.1). We follow standard practice by augmenting node features with positional encodings (PEs) [45], noting that our partial EVD provides PEs for free. Additional implementation details are provided in Appendix˜D.
Wavelet methods categorization. To fairly assess wavelet-based approaches, we distinguish between methods based on their theoretical adherence. Theory-compliant (TC) methods satisfy admissibility constraints and propagate entirely through wavelet operations. Only SGWT [22] and LR-GWN (with admissibility) qualify as theory-compliant. Theory-relaxed (TR) methods either incorporate auxiliary non-wavelet components (WaveGC [28]) or relax theoretical constraints (LR-GWN (without admissibility), GWNN [47], DEFT [5], ASWT-SGNN [30]).
@@ 5.1 Long-Range Tasks
Datasets. We evaluate on Peptides-Func (multi-label classification of 10 functional classes) and Peptides-Struct (node-level regression for 11 3D structural properties) from the Long-Range Graph Benchmark [15, 45]. These tasks specifically test long-range propagation in biological structures where distant interactions determine biological functions.
Results. LR-GWN achieves state-of-the-art performance on both datasets, as reported in Table˜1. Most notably, it significantly outperforms all existing theory-compliant methods (70.52 vs 60.23 on Peptides-Func) while maintaining theoretical guarantees. When constraints are relaxed, it surpasses all theory-relaxed baselines. It is worth noting that even the theory-compliant version of LR-GWN, achieves comparable performance to the theory-relaxed baselines, while maintaining pure wavelet interpretability and propagation.
Table 1: Benchmark results by theoretical adherence. Theory-compliant methods (TC) satisfy wavelet admissibility and use pure wavelet propagation. Theory-relaxed methods (TR) either use additional non-wavelet modules, or relax theory constraints. Bold/underline represent first and second best, respectively.
(a) Long-range benchmarks.
@@
SGWT [22] 60.23 $\pm$ 0.27 - 25.39 $\pm$ 0.21 -
GWNN [47] - 65.47 $\pm$ 0.48 - 27.34 $\pm$ 0.04
DEFT [5] - 66.95 $\pm$ 0.63 - 25.06 $\pm$ 0.13
WaveGC [28] - 69.73 $\pm$ 0.43 - 24.95 $\pm$ 0.07
@@
(b) Short-range benchmarks.
@@
SGWT [22] 92.45 $\pm$ 0.15 - 85.19 $\pm$ 0.22 -
GWNN [47] - 94.45 $\pm$ 0.18 - 90.75 $\pm$ 0.16
ASWT-SGNN [29] - 93.80 $\pm$ 0.12 - 89.40 $\pm$ 0.19
WaveGC [28] - 95.37 $\pm$ 0.44 - 92.26 $\pm$ 0.18
@@
5.2 Short-Range Tasks
Datasets. We evaluate on Photo and Computers from Amazon co-purchase graphs [36, 42]. These datasets focus on local neighborhood interactions, validating that our long-range optimization does not compromise local modeling capabilities.
Results. As reported in Table˜1, LR-GWN achieves state-of-the-art performance on Photo and competitive results on Computers, demonstrating our proposed hybrid parametrization offers the flexibility to model both global and local propagation. Contrary to long-range tasks, relaxing admissibility constraints does not consistently improve performance on local interaction datasets.
Training Time. On Photo, training LR-GWN takes approximately 3s per epoch with total training time around 20 minutes for the full run (400 epochs). The partial EVD preprocessing with $k=50$ costs roughly 1.3s using sparse solvers, accounting for approximately 0.1% of total training time.